Session: Jun 2007 (session 2)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Onsala (Michael Lindqvist at Wed Jun 13 14:40:56 2007) Success :
No known problems.

Medicina (Mauro at Wed Jun 13 14:59:03 2007) Success :
some clouds in yhan sky

Torun (K. Borkowski at Thu Jun 14 06:41:32 2007) Minor failures :
Lost 4-minute scan at 13:46 UTC due to a problem in the telescope azimuth drive.

Torun (K. Borkowski at Thu Jun 14 07:13:34 2007) Minor failures :
I forgot to enter an info important for the correlator: we recorded this experiment with swapped polarizations! The channels were reconnected correctly immediately after this experiment.

Effelsberg (A. Polatidis at Thu Jun 14 14:14:24 2007) Minor failures :
The focus mechanism was not active until 0500 UT, leading to reduced gain for the first 2.5 hours. Otherwise OK.

Noto (P. Cassaro at Mon Jun 18 13:15:01 2007) Success :